OpenStudy (anonymous):

2 square root of 45 end root minus 2 square root 5

OpenStudy (anonymous):

try mathway.com it's a calculator that will do any of these kind of problems. :)

OpenStudy (campbell_st):

well look at it like this you need to express 45 as the product of 2 factors, 1 of the factors being a square number. then use some of the laws for radicals. \[2\sqrt{45} = 2 \sqrt{9 \times 5} = 2\sqrt{9} \times 2 \sqrt{5} = 6\sqrt{5}\] so the problem is now \[6\sqrt{5} - 2 \sqrt{5} = \]